OpenStudy (anonymous):

3 - 5(x + 2) = 3(x + 1) + 6 explain steps please

OpenStudy (anonymous):

We're going to use PEMDAS. First, solve all the parentheses. Using the distributive rule, 5(x+2) becomes 5x+10, because 5*x=5x and 5*2=10. Doing the same thing to the parentheses on the other side, the equation becomes 3-5x+10=3x+3+6. Next, we combine all like terms. So on the left side, we combine +3 and +10 to get +13. On the right side, we combine +3 and +6 to get +9. So the equation becomes -5x+13=3x+9. Then, we make sure there is only one x. Thus, we add 5x to both sides giving us 13=8x+9. After that, we get the x all alone. To do this, we subtract 9 from both sides. The equation then becomes 8x=4. Finally, divide by 8 to get x=4/8 or x=1/2.

OpenStudy (anonymous):

Let me know if you need more help.

OpenStudy (anonymous):

thnks, I already got the answer. It's -2 but thnks anyway. Algebra is hard as hell.

OpenStudy (anonymous):

3 - 5(x + 2) = 3(x + 1) + 6 First I take what’s within the parenthesis of both sides and multiply them by 5(the first) and 3 (second). 3-5x -10 = 3x+ 3 +6 Next I have to combine like terms on both sides. -5x -7 = 3x +9 After that I divide 16 by -8 -8x=16 to x= 16/-8 And here’s my answer x=-2

OpenStudy (anonymous):

Wow, i completely missed the negative sign on my first distribution. *facepalm*

OpenStudy (anonymous):

dude, i'm sure i'm worse at math than you are. flvs

OpenStudy (anonymous):

this is one of my lucky days

OpenStudy (anonymous):

Yeah, but i always make that mistake.

OpenStudy (anonymous):

thnks though. Peace out.

OpenStudy (anonymous):

see ya